# An analogue of Ramanujan’s identity for Bernoulli-Carlitz numbers

### v3

@@ 1. Introduction
(1.2) $\frac{z}{e^{z}-1}=\sum_{m=0}^{\infty}B_{m}\frac{z^{m}}{m!}.$
Then Ramanujan discovered the following formula for the special values of $\zeta(s)$ at the odd positive integers (see Entry 21(i) in Ramanujan’s second notebook [3, p. 173] or his lost notebook [1, p. 319–320, formula (28)]).
@@
(1.3) $\alpha^{-m}$$\left\{\dfrac{1}{2}\,\zeta(2m+1)+\sum_{n=1}^{\infty}\dfrac{n^{-2m-1}}{e^{2\alpha n}-1}\right\}$
@@
In 1925, Marulkar [16] gave the first proof of (1.3), but he did not aware that this formula can be found in Ramanujan’s Notebooks. Then (1.3) has been rediscovered and studied by Grosswald in [8, 9]. In 1977, Berndt [2, Theorem 2.2] showed that both Euler’s formula (1.1) and Ramanujan’s formula (1.3) are special cases of a general transformation formula of analytic Eisenstein series, thus they are natural companions of each other. And in 2011, Gun, Murty and Rath [14] gave a modern interpretation of (1.3), that is, it encodes fundamental transformation properties of Eisenstein series on the full modular group and their Eichler integrals. This approach has been further extended by Berndt and Straub [4, Section 5] to weight $12k+$ 1 Eisenstein series of level 2 through secant Dirichlet series. Recently, Chavan [7] presented an elementary proof of (1.3), which relies on a Mittag-Leffler type expansion of hyperbolic contangent function
(1.4) $\coth\left(\pi x\right)=\dfrac{1}{\pi x}+\dfrac{2x}{\pi}\sum_{k=1}^{\infty}\dfrac{1}{x^{2}+k^{2}}$
for $x\in\mathbb{C}\setminus{\{0\}}$ and Euler’s identity for even zeta values (1.1).
In this note, we shall investigate the above Ramanujan’s identity in the functions fields setting. In this situation, an analogy of the Riemann zeta function $\zeta(s)$ was introduced by Goss [12] in 1979. Let $q=p^{k}$ be a power of a prime number $p$, $\mathbb{F}_{q}$ be the finite field with $q$ elements and $\mathbb{F}_{q}^{*}$ be its multiplicative group. Let $R=\mathbb{F}_{q}[T]$ be the polynomial ring with one variable over the finite field $\mathbb{F}_{q}$ and $K=\mathbb{F}_{q}(T)$ be the rational function field. Let $K_{\infty}=\mathbb{F}_{q}((\frac{1}{T}))$ be the completion of $K$ at the infinite place $\infty=\left(\frac{1}{T}\right)$, $K_{\infty}^{*}=K_{\infty}\setminus\{0\}$ be the multiplicative group of the field $K_{\infty}$ and $\mathbb{C}_{\infty}$ be the completion of an algebraic closure of $K_{\infty}$. Let $R_{1}$ be the set of monic polynomials in $R,$ which is an analogue of the set of positive integers. Goss considered the infinite series
(1.5) $\zeta_{\infty}(s)=\sum_{A\in R_{1}}A^{-s}$
@@
$[j]=T^{q^{j}}-T,\quad D_{j}=[j]{[j-1]}^{q}\cdots{[1]}^{q^{j-1}}$
with $D_{0}=1$. The Carlitz exponential is defined by
(1.6) $e(z)=\sum_{j=0}^{\infty}\frac{z^{q^{j}}}{D_{j}}.$
@@
(1.7) $e(z)=e_{\tilde{\pi}R}(z)=z\prod_{0\neq\lambda\in\tilde{\pi}R}\left(1-\frac{z}{\lambda}\right),$
where $\tilde{\pi}\in\mathbb{C}_{\infty}$ is an analogy of the classical $2\pi i~{}(i=\sqrt{-1})$, as it is a fundamental period of Carlitz’s exponential $e(z)$. In 1941, Wade [19] proved that $\tilde{\pi}$ is transcendental over $K$.
Let $\Omega=\mathbb{C}_{\infty}\setminus K_{\infty}$ be Drinfeld’s upper half-plane, which is an analogue of the classical upper half-plane $H=\{\tau\in\mathbb{C}~{}|~{}\textrm{Im}(\tau)>0\}.$ Recently, by applying $e(z)$, Hamahata [15] introduced a function fields analogue of Carlitz’s generalized Dedekind sums,
(1.8) $t_{r}^{(n)}(a,c)=\sum_{0\neq\mu\in R/cR}e(\tilde{\pi}\mu/c)^{-n-1+r}e(\tilde{\pi}a\mu/c)^{-r}$
for $a,c\in R\setminus\{0\}$ and $r\in\{0,1,\ldots,n+1\}$. It enables him to establish the Apostol–type transformation formulas for the following Lambert series in function fields,
(1.9) $\xi_{n}(z)=\sum_{a\in R_{1}}\frac{1}{a^{n}}\frac{1}{e(\tilde{\pi}az)},~{}~{}z\in\Omega.$
(See [15, Section 4]).
@@
(1.10) $\frac{z}{e(z)}=\sum_{m=0}^{\infty}\frac{BC_{m}}{\Gamma_{m}}{z^{m}}.$
As in the classical case, the Bernoulli-Carlitz numbers have many deep applications in the arithmetic of function fields, such as Kummer’s and Herbrand’s criterions for the class groups of cyclotomic function fields (see [13, Section 9.2], [18, Section 5.2] or [10, 11]). In 1930s, Carlitz [5, 6] showed an analogue of Euler’s formula (1.1), that is, the special values of $\zeta_{\infty}(s)$ at the “even” positive integers $m=(q-1)j~{}(j\in\mathbb{N})$ are given by
(1.11) $\zeta_{\infty}(m)=\tilde{\pi}^{m}\frac{BC_{m}}{\Gamma_{m}}.$
@@
If $\alpha$ and $\beta$ are elements in $\mathbb{C}_{\infty}$ such that $\alpha\beta=\tilde{\pi}^{2}$ and if $j$ is a positive integer, then we have
$\alpha^{-(\frac{q-1}{2})(j-1)-(q-2)}$ $\sum_{A\in R_{1}}\frac{A^{-(q-1)j-(q-2)}}{e(\alpha A)}+\beta^{-(\frac{q-1}{2})(j-1)-(q-2)}\sum_{A\in R_{1}}\frac{A^{-(q-1)j-(q-2)}}{e(\beta A)}$
(1.12) $=\sum_{k=0}^{j+1}\frac{BC_{(q-1)k}}{\Gamma_{(q-1)k}}\frac{BC_{(q-1)(j-k+1)}}{\Gamma_{(q-1)(j-k+1)}}\alpha^{(\frac{q-1}{2})k}\beta^{(\frac{q-1}{2})(j-k+1)}.$
@@
$\sum_{A\in R_{1}}\frac{A^{-(q-1)j-(q-2)}}{e(\alpha A)}\quad\text{and}\quad\sum_{A\in R_{1}}\frac{A^{-(q-1)j-(q-2)}}{e(\beta A)}$

### v4

@@ 1. Introduction
$\frac{z}{e^{z}-1}=\sum_{m=0}^{\infty}B_{m}\frac{z^{m}}{m!}.$ (1.2)
Then Ramanujan discovered the following formula for the special values of $\zeta(s)$ at the odd positive integers (see Entry 21(i) in Ramanujan’s second notebook [4, p. 173] or his lost notebook [2, p. 319–320, formula (28)]).
@@
$\alpha^{-m}$ $\left\{\dfrac{1}{2}\,\zeta(2m+1)+\sum_{n=1}^{\infty}\dfrac{n^{-2m-1}}{e^{2\alpha n}-1}\right\}$ (1.3)
@@
In 1925, Marulkar [24] provided the first proof of equation (1.3), though he was not aware that this formula had already appeared in Ramanujan’s notebooks. Later, (1.3) was rediscovered and studied by Grosswald in [11, 12]. In 1977, Berndt [3, Theorem 2.2] demonstrated that both Euler’s formula (1.1) and Ramanujan’s formula (1.3) are special cases of a general transformation formula for analytic Eisenstein series, establishing them as natural companions. In 2011, Gun, Murty, and Rath [21] offered a modern interpretation of (1.3), showing that it encodes fundamental transformation properties of Eisenstein series on the full modular group and their Eichler integrals. This approach was further extended by Berndt and Straub [5, Section 5] to weight $2k+1$ Eisenstein series of level 2 via secant Dirichlet series. In 2024, as a complement to the work of Berndt and Straub [5], Dixit [10] provided a survey of recent developments and generalizations of Ramanujan’s identity (1.3), covering extensions in settings such as Lambert series, Koshliakov zeta functions, higher Herglotz functions, and non-holomorphic Eisenstein series. The survey also discussed connections to modular forms, number theory, and other mathematical areas that have emerged in the last decade.
Recently, Chavan [8] presented an elementary proof of (1.3), based on a Mittag-Leffler type expansion of the hyperbolic cotangent function:
$\coth(\pi x)=\frac{1}{\pi x}+\frac{2x}{\pi}\sum_{k=1}^{\infty}\frac{1}{x^{2}+k^{2}},$ (1.4)
for $x\in\mathbb{C}\setminus\{0\}$, and Euler’s identity for even zeta values (1.1).
In [9], the authors generalized Ramanujan’s identity (1.3) and situated it within a broader analytic framework. Specifically, Ramanujan’s identity is derived as a special case of their master identity (Theorem 2.5) when a=2m1a=-2m-1, this master identity provides a unified transformation for the Lambert series $\sum_{n=1}^{\infty}\sigma_{a}(n)e^{-ny}$, covering both modular and non-modular cases, and for an odd integer $a$, the master identity reduces to classical modular transformations (e.g., Eisenstein series on $\mathrm{SL}_{2}(\mathbb{Z})$), which are connected to Ramanujan’s formula. Furthermore, the authors also derived a new companion formula (Corollary 2.13) involving even zeta values $\zeta(2m)$, contrasting with Ramanujan’s result for odd zeta values. Thus, the paper not only recovers Ramanujan’s result but also extends it to a broader class of series and provides new identities in the same spirit.
In this note, we shall investigate the above Ramanujan’s identity in the functions fields setting. In this situation, an analogy of the Riemann zeta function $\zeta(s)$ was introduced by Goss [16] in 1979. Let $q=p^{k}$ be a power of a prime number $p$, $\mathbb{F}_{q}$ be the finite field with $q$ elements and $\mathbb{F}_{q}^{*}$ be its multiplicative group. Let $A=\mathbb{F}_{q}[T]$ be the polynomial ring with one variable over the finite field $\mathbb{F}_{q}$ and $K=\mathbb{F}_{q}(T)$ be the rational function field. Let $K_{\infty}=\mathbb{F}_{q}((\frac{1}{T}))$ be the completion of $K$ at the infinite place $\infty=\left(\frac{1}{T}\right)$, $K_{\infty}^{*}=K_{\infty}\setminus\{0\}$ be the multiplicative group of the field $K_{\infty}$ and $\mathbb{C}_{\infty}$ be the completion of an algebraic closure of $K_{\infty}$. Let $A_{+}$ be the set of monic polynomials in $A,$ which is an analogue of the set of positive integers. Goss considered the infinite series
$\zeta_{\infty}(s)=\sum_{a\in A_{+}}a^{-s}$ (1.5)
@@
$j]=T^{q^{j}}-T,\quad D_{j}=[j]{[j-1]}^{q}\cdots{[1]}^{q^{j-1}}$
with $D_{0}=1$. Let $C$ be the Carlitz module and the Carlitz exponential $e_{C}(z)$ associated with $C$ is defined by
$e_{C}(z)=\sum_{j=0}^{\infty}\frac{z^{q^{j}}}{D_{j}}.$ (1.6)
@@
$e_{C}(z)=e_{\tilde{\pi}A}(z)=z\prod_{0\neq\lambda\in\tilde{\pi}A}\left(1-\frac{z}{\lambda}\right),$ (1.7)
where $\tilde{\pi}\in\mathbb{C}_{\infty}$ is an analogy of the classical $2\pi i~(i=\sqrt{-1})$, as it is a fundamental period of Carlitz’s exponential $e_{C}(z)$. In 1941, Wade [29] proved that $\tilde{\pi}$ is transcendental over $K$.
@@
$\frac{z}{e_{C}(z)}=\sum_{m=0}^{\infty}\frac{BC_{m}}{\Gamma_{m}}{z^{m}}.$ (1.9)
As in the classical case, the Bernoulli-Carlitz numbers have many deep applications in the arithmetic of function fields, such as Kummer’s and Herbrand’s criterions for the class groups of cyclotomic function fields (see [19, Section 9.2], [28, Section 5.2] or [13, 15]). In 1930s, Carlitz [6, 7] showed an analogue of Euler’s formula (1.1), that is, the special values of $\zeta_{\infty}(s)$ at the “even” positive integers $m=(q-1)j~(j\in\mathbb{N})$ are given by
$\zeta_{\infty}(m)=\tilde{\pi}^{m}\frac{BC_{m}}{\Gamma_{m}}.$ (1.10)
@@
Let $e_{C}(z)$ be Carlitz’s exponential function (1.6), $\Gamma_{m}$ be Carlitz’s factorial (1.8) and $BC_{m}$ be the Bernoulli-Carlitz numbers (1.9). If $\alpha$ and $\beta$ are elements in $\mathbb{C}_{\infty}$ such that $\alpha\beta=\tilde{\pi}^{2}$ and if $j$ is a positive integer, then we have
$\alpha^{-(\frac{q-1}{2})(j-1)-(q-2)}$ $\sum_{a\in A_{+}}\frac{a^{-(q-1)j-(q-2)}}{e_{C}(\alpha a)}+\beta^{-(\frac{q-1}{2})(j-1)-(q-2)}\sum_{a\in A_{+}}\frac{a^{-(q-1)j-(q-2)}}{e_{C}(\beta a)}$
$=\sum_{k=0}^{j+1}\frac{BC_{(q-1)k}}{\Gamma_{(q-1)k}}\frac{BC_{(q-1)(j-k+1)}}{\Gamma_{(q-1)(j-k+1)}}\alpha^{(\frac{q-1}{2})k}\beta^{(\frac{q-1}{2})(j-k+1)}.$ (1.11)
@@
$\sum_{a\in A_{+}}\frac{a^{-(q-1)j-(q-2)}}{e_{C}(\alpha a)}\quad\text{and}\quad\sum_{a\in A_{+}}\frac{a^{-(q-1)j-(q-2)}}{e_{C}(\beta a)}$
@@
The subsequent sections of this paper are organized as follows. In Section 2, we provide some remarks that connect the above result with the theory of modular forms over function fields, including relationships between Petrov’s $A$-expansions and Hamahata’s construction of Lambert series. In Section 3, we present the proof of Theorem 1.2.
2. On its relations between modular forms in function fields
In this section, as suggested by the referee, we give some remarks which relate Theorem 1.2 with the theory of modular forms in the function fields.
2.1. Petrov’s $A$-expansions
In the classical setting, let the upper half-plane be $\mathbb{H}=\{\tau\in\mathbb{C}\mid\operatorname{Im}(\tau)>0\}$. The modular group $\mathrm{SL}_{2}(\mathbb{Z})$ acts on $\mathbb{H}$ by fractional linear transformations: for a matrix $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\mathrm{SL}_{2}(\mathbb{Z})$, the action is defined by $\tau\mapsto\frac{a\tau+b}{c\tau+d}$. For any odd integer $n\geq 1$, recall that the Lambert series is defined as
$G_{n}(x)=\sum_{m=1}^{\infty}\frac{m^{-n}x^{m}}{1-x^{m}},\quad|x|<1.$ (2.1)
By using $G_{n}(x)$, Ramanujan’s identity (1.3) can be wrtten as
$\alpha^{-m}$ $\left\{\dfrac{1}{2}\,\zeta(2m+1)+G_{2m+1}\left(e^{-2\alpha}\right)\right\}$ (2.2)
$\quad\quad\quad\quad\quad\quad\quad-(-\beta)^{-m}\,\left\{\dfrac{1}{2}\,\zeta(2m+1)+G_{2m+1}\left(e^{-2\beta}\right)\right\}$
Let
$\mathcal{G}_{2k}(z):=\sum_{(0,0)\neq(m,n)}\frac{1}{(mz+n)^{2k}}$
and
$\mathcal{E}_{2k}(z):=\mathcal{G}_{2k}(z)/2\zeta(2k)$
be the Eisenstein series and its normalized version, respectively. From the Fourier expansion of $\mathcal{G}_{2k}(z)$, we have the following Lambert expansion for $\mathcal{E}_{2k}(z)$:
$\mathcal{E}_{2k}(z)=1+\frac{2}{\zeta(1-2k)}\sum_{n=1}^{\infty}n^{2k-1}\frac{q^{n}}{1-q^{n}}.$
Denote by $q_{n}:=q^{n}$ and G(x):=x/(1x)G(x):=x/(1-x). Applying (2.1), we may rewrite the above expansion as
$\mathcal{E}_{2k}(z)=1+\frac{2}{\zeta(1-2k)}\sum_{n=1}^{\infty}n^{2k-1}G(q_{n}).$ (2.3)
Now we come to the situation of function fields. Let $\Omega=\mathbb{C}_{\infty}\setminus K_{\infty}$ be Drinfeld’s upper half-plane, which is an analogue of the classical upper half-plane $\mathbb{H}=\{\tau\in\mathbb{C}~|~\textrm{Im}(\tau)>0\}.$ Let $\Gamma=\rm{GL}_{2}(A)$, $k$ be a positive integer and $m$ be an integer with $0\leq m<q-1$. In the function fields setting, in analogy with the classical $\rm{SL}_{2}(\mathbb{Z})$-theory, Goss and Gekeler introduced the modular forms of weight $k$ and type $m$ on $\Omega$ associated with $\Gamma$, named the Drinfeld modular forms (see [17, 18, 14]). The space of such forms is denoted by $M_{k,m}=M_{k,m}(\Gamma)$, and the subspaces of cusp forms is denoted by $S_{k,m}=S_{k,m}(\Gamma)$.
Let
$u:=u(z)=\frac{1}{e_{C}(\tilde{\pi}z)}.$ (2.4)
It is regular on $\Omega$ and is the usual uniformizer at $\infty$. For $a\in A_{+}$ of degree $d$, let
$u_{a}:=u(az)=\frac{1}{e_{C}(\tilde{\pi}az)}.$ (2.5)
In function fields, similarly, we define the Eisenstein series of weight $k$ with $k\equiv 0\pmod{q-1}$ by
$\mathcal{G}_{k}(z):=\sum\frac{1}{(az+b)^{k}},$
where the sum is taken over all nonzero pairs $(a,b)\in A\times A$. As shown in [14], this function admits a decomposition involving its constant term and an $A$-expansion. Specifically, we have
$\frac{\mathcal{G}_{k}(z)}{\tilde{\pi}^{k}}=\sum_{a\in A_{+}}(\tilde{\pi}a)^{-k}+\sum_{a\in A_{+}}G_{k,\Lambda_{C}}(u_{a}),$ (2.6)
where $G_{k,\Lambda_{C}}(z)$ denotes the Goss polynomial associate with the Carlitz module $C$ (see [20, p. 332, Definition 1]).
This identity illustrates that the Eisenstein series $\mathcal{G}_{k}(z)$ can be expressed as a sum of a constant term (a “zeta-value” at a positive integer) and a series over monic polynomials $a\in A_{+}$, each term involving the function $G_{k,\Lambda_{C}}(u_{a})$. This structure serves as a prototype for the more general $A$-expansions introduced by Petrov.
In [25], Petrov built on the above example and constructed families of cusp forms with $A$-expansions.
Theorem 2.1 (Petrov, see [25, Theorem 1.3] or [20, Theorem 1]).
Let $k,n$ be two positive integers such that k2nk-2n is a positive multiple of q1q-1 and $n\leq p^{\operatorname{ord}_{p}(k-n)}$. Then
$f_{k,n}:=\sum_{a\in A_{+}}a^{k-n}G_{n,\Lambda_{C}}(u_{a})$ (2.7)
is an element of $S_{k,m}$ with $n\equiv m\pmod{q-1}$.
The above construction led Goss to introduce $\mathfrak{v}$-adic modular forms in function fields in the sense of Serre (see [20, p. 334, Definition 4]), while Serre [26] started to define the $p$-adic modular forms in number fields as $p$-adic $q$-expansion which are uniform limits of $q$-expansions of the classical Eisenstein series.
Although similar expansions of Eisenstein series exist in the function field setting due to Gekeler and Petrov (see equations (2.6) and (2.7)), the situation is quite different from that in number fields. This is because the corresponding term $G_{n,\Lambda_{C}}(u_{a})$ does not possess a transformation formula analogous to that of the classical Lambert series (2.9) (see [23]). For the same reason, the transformation formula of Eisenstein series and its expansions (2.6) may not be used to derive the main result (Theorem 1.2).
In fact, the transformation formulae for the logarithm of the Dedekind eta function in number fields and for the Dedekind sum in function fields were discussed in [22, 23]. These works also define a new Lambert series in function fields as an analogue of these transformation laws (see (2.10)).
Furthermore, as remarked by Goss [20]: “in a certain sense, Petrov’s construction of cusp forms via $A$-expansions is both analogous to and ‘orthogonal to’ (i.e., it lies in the space of cusp forms, rather than Eisenstein series) the classical construction of Eisenstein series” (see [20, p. 334, Remark 3]). In the sense of this analogy, comparing (2.3) and (2.7), we observe that the term $\zeta(1-2k)$ appears in the Lambert expansions of the classical Eisenstein series, whereas the $A$-expansion (2.7) does not contain any term analogous to $\zeta(1-2k)$. Therefore, we believe it is reasonable that although the term $\zeta(2m+1)$ appears in the classical Ramanujan’s identity (1.3), it does not appear in its function field analogue (1.11).
2.2. Hamahata’s Lambert series
In the classical setting, for any $\tau\in\mathbb{H}=\{\tau\in\mathbb{C}\mid\operatorname{Im}(\tau)>0\}$, the Dedekind eta function is defined as
$\eta(\tau)=e^{\pi i\tau/12}\prod_{n=1}^{\infty}\left(1-e^{2\pi in\tau}\right).$
Dedekind proved the transformation formula for $\log\eta(\tau)$, that is, for a matrix $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\mathrm{SL}_{2}(\mathbb{Z})$ with $a\neq 0$ and $c>0$, we have
$\log\eta\left(\frac{az+b}{cz+d}\right)=\log\eta(z)+\frac{1}{2}\log\left(\frac{cz+d}{i}\right)+\frac{\pi i(a+d)}{12c}-\pi iD(a,c),$ (2.8)
where $D(a,c)$ is the classical Dedekind sum, defined for coprime integers $a$ and $c>0$ by
$D(a,c)=\frac{1}{4c}\sum_{k=1}^{c-1}\cot\left(\frac{\pi ak}{c}\right)\cot\left(\frac{\pi k}{c}\right).$
Since by (2.1) we have
$\log\eta(\tau)=\frac{\pi i\tau}{12}-G_{1}(e^{2\pi i\tau}),$
the transformation formula for the Lambert series $G_{1}(e^{2\pi i\tau})$ can be derived from (2.8).
For the general Lambert series $G_{n}(e^{2\pi i\tau})$ ($n>1$), in 1950 Apostol [1] proved their transformation formulas, that is, for odd integer $n>1$ and $\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\mathrm{SL}_{2}(\mathbb{Z})$, we have
$(c\tau+d)^{n-1}G_{n}\left(e^{2\pi i\gamma\tau}\right)$ $=G_{n}\left(e^{2\pi i\tau}\right)+\frac{1}{2}\zeta(n)\left\{1-(c\tau+d)^{n-1}\right\}$ (2.9)
$\quad-\frac{(2\pi i)^{n}}{2(n+1)!}\sum_{r=0}^{n+1}\binom{n+1}{r}(c\tau+d)^{n-r}D_{r}(-a,c),$
where
$D_{r}(a,c)=\sum_{\mu=1}^{c-1}\overline{B}_{n+1-r}\left(\frac{a\mu}{c}\right)\overline{B}_{r}\left(\frac{\mu}{c}\right)$
is the generalized Dedekind sum and $\overline{B}_{n}(x)$ be the periodic Bernoulli functions.
In the function fields situation, as remarked in the previous subsection, although the $A$-expansion (2.6) and (2.7) can be viewed as an analogy of the Fourier expansion for Eisenstein series (2.3), it is not reasonable to view the corresponding terms $G_{n,\Lambda_{C}}(u_{a})$ as an analogy of Lambert series, because it does not have a transformation formula similar to (2.9). To fill this gap, Hamahata [23, Definition 2.1] introduced the following new series $\xi_{n}(z)$ in the function fields setting.
Definition 2.2 (Hamahata, see [23, p. 275, Definition 2.1]).
For a natural number $n\in\mathbb{N}$, let
$\xi_{n}(z)=\sum_{a\in A_{+}}\frac{1}{a^{n}}u(az).$ (2.10)
And he also proved an analogy of Apostol’s transformation formula (2.9):
Theorem 2.3 (Hamahata, see [23, p. 281, Theorem 4.4]).
Assume that $q-1\mid n+1$. For $\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in GL_{2}(A)$ with $a\neq 0$, $c\neq 0$, we have
$(cz+d)^{n-1}\xi_{n}(\gamma z)=(\det\gamma)^{-1}\Bigg[\xi_{n}(z)+\frac{\alpha(n+1)}{\pi c^{n}}\left\{(cz+d)^{n}+\frac{1}{cz+d}\right\}\\ -\frac{1}{\pi c^{n}}\sum_{r=1}^{n}\alpha(r)\alpha(n+1-r)(cz+d)^{r-1}\\ -\frac{\pi^{n}}{c^{n}}\sum_{r=1}^{n}(\det\gamma)^{r-1}(cz+d)^{n-r}t_{r}^{(n)}(a,c)\Bigg].$ (2.11)
Here $t_{r}^{(n)}(a,c)$ denotes the function fields analogue of Carlitz’s generalized Dedekind sums:
$t_{r}^{(n)}(a,c)=\sum_{0\neq\mu\in A/cA}e_{C}(\tilde{\pi}\mu/c)^{-n-1+r}e_{C}(\tilde{\pi}a\mu/c)^{-r}$ (2.12)
for $a,c\in A\setminus\{0\}$ and $r\in\{0,1,\ldots,n+1\}$. Combining (2.4), (2.5) and (2.10), by noticing that $\alpha\beta=\tilde{\pi}^{2}$, we can state Theorem 1.2 in the form of Hamahata’s Lambert series $\xi_{n}(z)$, which is an analogy of the classical formula (2.2) above.
Corollary 2.4.
Let $\xi_{n}(z)$ be Hamahata’s Lambert series (2.10), $\Gamma_{m}$ be Carlitz’s factorial (1.8) and $BC_{m}$ be the Bernoulli-Carlitz numbers (1.9). If $\alpha$ and $\beta$ are elements in $\mathbb{C}_{\infty}$ such that $\alpha\beta=\tilde{\pi}^{2}$ and if $j$ is a positive integer, then we have
[…]
$=\sum_{k=0}^{j+1}\frac{BC_{(q-1)k}}{\Gamma_{(q-1)k}}\frac{BC_{(q-1)(j-k+1)}}{\Gamma_{(q-1)(j-k+1)}}\alpha^{(\frac{q-1}{2})k}\beta^{(\frac{q-1}{2})(j-k+1)}.$ (2.13)
As demonstrated in [9], our result also opens the possibility of placing Ramanujan’s identity within a broader analytic framework in the function field setting.